# Always Keep Your Promises: DynamicLRP, A Model-Agnostic Solution To Layer-Wise Relevance Propagation

### v1

@@ 6.1 Architecture-Specific Factors on LRP Performance
Table 4: Architecture-specific values relating to the Promise System’s impact on LRP efficiency. We include a separate row for the (F)ine(T)uned version of Flan-T5-large which uses LoRA adapters, changing some of these values drastically. Depth is in terms of either number of (T)ransformer layers or (C)onvolutional layers.
@@ 7 Discussion
The recorded values from Table 4 provide insight on how the Promise System affects runtime and provide empirical $|V_{P}|$ and $\delta$ to help reason about the theoretical memory and runtime complexities. The most striking observation comes from the difference in $\rho$ between the two versions of Flan-T5-large. While the finetuned T5 has nearly twice as many autograd nodes and much higher $\delta$, the number of Promises does not increase by nearly the same scale.
The memory improvement of AttnLRP over our LRP is likely due to the fact that it does not require the full autograd computation graph to be stored, only the subgraphs associated with the above mentioned layers/modules, as well as activations and normalization layers. This is characteristic to the implementation, which is built on the Zennit library Anders et al. (2021). Any additional operations performed do not have autograd nodes saved, and thus reduce total peak VRAM usage, as opposed to our method which tracks such operations.
7.1 Operation Coverage
By breaking LRP down to the level of tensor operations, we knowingly take on the task of covering a sufficiently large portion of the operation space to make this useful across a wide range of models. However, consider that the set of tensor operations we are interested in, specifically mathematical and shape operations, is a bounded and well-defined set. This places a tangible and conceptual limit on our implementation scope.
There are also case-by-case implementation considerations. In shape operations, which manipulate only the structure of tensors, the task of relevance propagation reduces to accumulating relevance at the inputs which map to the output. Here, we simply reuse the functions from gradient backpropagation, which do exactly this and are produced by the auto-differentiation library. Activation functions such as ReLU, GELU, etc. have been shown to correspond to Identity propagation functions, requiring no implementation overhead. Finally, fused/composite operations can re-routed or decomposed during auxiliary graph construction (Algorithm 1) into components we already handle.
Table 5: DynamicLRP operation coverage breakdown of 15 models spanning vision, text, and audio modalities. Unique Ops indicates the number of distinct operation types in each model’s computation graph.
Model Modality Covered Nodes Unique Ops covered
VGG16 Simonyan & Zisserman (2015) Vision 90/90 9/9
ResNet-50 He et al. (2015) Vision 339/339 10/10
ViT-b-16 Dosovitskiy et al. (2021) Vision 779/779 16/16
EfficientNetv2-medium Tan & Le (2021) Vision 1,526/1,526 12/12
Wav2Vec2-xls-r-300m Babu et al. (2021) Audio 2,021/2,022 18/19
whisper-small Radford et al. (2022) Audio 2,729/2,729 22/22
GPT-2 Radford et al. (2019) Language 885/885 19/19
RoBERTa-large Liu et al. (2019) Language 1,461/1,461 14/14
Llama3.2-1B Dubey & et al. (2024) Language 1,787/1,787 24/24
Gemma-3-270m-it Team (2025a) Language 2,482/2,482 24/24
Qwen3-0.6B Team (2025b) Language 2,590/2,590 18/18
Flan-T5-large Chung et al. (2022) Language 5,713/5,713 24/24
DePlot Liu et al. (2022) Vision/Language 2,863/2,863 23/23
Mamba-130m Gu & Dao (2024) State Space 3,397/3,421 25/26
Aggregate - 31,440/31,465 -
Table 6: Uncovered operation breakdown, where applicable.
Model Uncovered Ops Node count
Wav2Vec2-xls-r-300m WeightNormInterface 1
Mamba-130m SplitWithSize 24
We validated model-agnostic coverage by analyzing computation graphs from 15 diverse architectures spanning vision (CNNs, ViTs), language (GPT, BERT, T5, Llama), audio (Wav2Vec2, Whisper), multimodal (DePlot), and state-space models (Mamba). As shown in Table 6, our implementation achieved 99.92% node coverage (31,440/31,465 nodes) across all architectures without any model-specific code modifications. The 25 uncovered nodes (0.08%) correspond to two operations: WeightNormInterface (weight normalization utility, 1 node) and SplitWithSize (variant of Split, 24 nodes). Both could be trivially supported by extending existing rules. We see that despite increasing model sizes and diversity, modern architectures are composed of only a small set of operations.
7.2 Extensibility to Other Deep Learning Frameworks
While it may seem that this design is strictly PyTorch-specific, we claim that the Promise-based Dynamic LRP can be extended to any deep learning framework with an auto-differentiation library.
We consider that differentiation is first and foremost a mathematical concept, and auto-differentiation stricly relies on computation graphs. This implies two important invariants across any auto-differentiation engine. Being built on computation graphs, these engines themselves are defined at the operation level, and for any function to be differentiated, the set of givens required is purely characteristic to the mathematical properties of the function in question, not the framework or programming language being used.
Thus, all implementations of auto-differentiation will be required to store the same minimum set of data for each function they share, in the structure of a computation graph. This signifies to us that the concept of operation-level LRP is sound and viable in any such deep-learning framework.
The Promise System allows even further flexibility to this, as its core purpose is the retrieval of forward pass activations not found in the computation graph during the backward pass. Therefore, any differences between frameworks in the backward context caching will also be overcome by our framework.
Appendix A List of Covered Operations
Below is the list of currently supported operations for DynamicLRP.
Category Operations
Arithmetic Add, Sub, Mul, Div, Neg, Sqrt, Rsqrt, Pow, Exp, Log
Linear Algebra Mm, Bmm, Convolution
Aggregation Sum, Mean, Cat, Stack, Unbind, Split, Max, Gather
Indexing and Selection Slice, Index, Select
Pooling MaxPool2D, MaxPool3D, AdaptiveAvgPool2D
Normalization NativeLayerNorm, NativeBatchNorm
Attention SDPA
Activations GELU, SiLU, ReLU, Softmax
Other MaskedFill, Clone, ToCopy, NativeDropout, Embedding
@@ Appendix B Algorithms
Algorithm 1 Computation Graph Construction
1:Model output hidden_stateshidden\_states
2:In-adjacency list in_adjin\_adj, Out-adjacency list out_adjout\_adj, Topologically sorted nodes topo_stacktopo\_stack
3:Initialize $in\_adj\leftarrow\emptyset$, $out\_adj\leftarrow\emptyset$
4:Initialize $visited\leftarrow\emptyset$, $topo\_stack\leftarrow\emptyset$
@@
6:DFS(rootroot, in_adjin\_adj, out_adjout\_adj, visitedvisited, topo_stacktopo\_stack) return in_adjin\_adj, out_adjout\_adj, topo_stacktopo\_stack
7:function DFS(fcnfcn, in_adjin\_adj, out_adjout\_adj, visitedvisited, topo_stacktopo\_stack)
8:  if $fcn=\text{None}$ or $fcn\in visited$ then
9:   return
10:  end if
11:  $visited\leftarrow visited\cup\{fcn\}$
12:  for each $child\in fcn.next\_functions$ do
13:   out_adj[fcn].append(child)out\_adj[fcn].append(child)
14:   in_adj[child].append(fcn)in\_adj[child].append(fcn)
15:   DFS(childchild, in_adjin\_adj, out_adjout\_adj, visitedvisited, topo_stacktopo\_stack)
16:  end for
17:  topo_stack.push(fcn)topo\_stack.push(fcn)
18:end function
Algorithm 2 Operation-Level Relevance Propagation With Promises
1:Model output hidden_stateshidden\_states, Relevance interpretation target target_nodetarget\_node, Auxiliary graph in_adjin\_adj, out_adjout\_adj, Output relevance $R_{out}$
2:Input relevance $R_{in}$
3:Initialize as in Algorithm 2
4:while stackstack do
5:  $curnode\leftarrow stack.pop()$
6:  curnode_in_rel=node_inputs[curnode]curnode\_in\_rel=node\_inputs[curnode]
7:  if curnode==target_nodecurnode==target\_node then return $R_{in}:=curnode\_in\_rel$
8:  else if curnodecurnode requires Promise then
9:   $curnode.promise\leftarrow create\_new\_promise(type(curnode),R_{out})$
10:   $curnode\_outputs\leftarrow curnode.promise.branches$
11:  else
12:   $prop\_fcn\leftarrow fcn\_map[type(curnode)]$
13:   $curnode\_outputs\leftarrow prop\_fcn(curnode\_in\_rel)$
14:  end if
15:  for each $child,output\in zip(out\_adj[curnode],curnode\_outputs)$ do
16:   node_inputs[child].add(output)node\_inputs[child].add(output)
17:   $nodes\_pending[child]\leftarrow nodes\_pending[child]-1$
18:   if nodes_pending[child]==0nodes\_pending[child]==0 then
19:     $stack.push(child)$
20:   end if
21:  end for
22:end while
@@
1:Autograd Node nodenode, Propagation input $R_{out}$
2:Propagation output list $R_{in}$
3:if $R_{out}$ is a Promise Branch then
4:  Define fwdfwd as a closure of nodenode’s forward pass.
5:  Define bwdbwd as a closure of nodenode’s relevance distribution logic.
6:  $R_{out}.record(fwd,bwd)$ return $R_{out}$
7:end if
8:// Propagate $R_{out}$ return $R_{in}$
@@
1:Autograd Node nodenode, Propagation input $R_{out}$
2:Propagation output list $R_{in}$
3:if $R_{out}$ is a Promise Branch then
4:  Define retrieve_fwd_outputretrieve\_fwd\_output as a function that extracts nodenode’s forward pass output given its saved tensors.
5:  activation=retrieve_fwd_output(node)activation=retrieve\_fwd\_output(node)
6:  $R_{out}.set\_arg(activation)$
7:  $R_{out}.trigger\_promise\_completion()$
8:  if $R_{out}.promise\_is\_complete$ then $R_{out}=R_{out}.propagated\_relevance$
9:  else
10:   // Signal for queueing this Node until the Promise is complete return
11:  end if
12:end if
13:// Propagate $R_{out}$ return $R_{in}$
Appendix C Implementation Details
Our implementation uses PyTorch hooks to record operations during the forward pass and constructs the computation graph dynamically. Each operation creates a promise object, which stores the propagation rule and tracks readiness and completeness.
Python Example: Promise System Below is a simplified Python pseudocode that captures the core logic of the Promise System:
class PromiseBranch:
  def __init__(self, promise_obj, index):
    self.promise = promise_obj
    self.promise_dict = promise.metadata_dict
    self.index = index
    self.fwd_chain = []
    self.bwd_chain = []
    self.children = [] # Recall that the child connection is from Branch -> Promise
  def record_fwd(self, fwd_closure):
    self.fwd_chain.append(fwd_closure)
  def record_bwd(self, bwd_closure):
    self.bwd_chain.append(bwd_closure)
  def setarg(self, val):
    # Brings args from the Arg Node to the Origin Node
    # Assumes that fwd_chain has recorded all node fwd closures leading up to an Arg Node
    for fcn in self.fwd_chain:
      # Iteratively apply the forward chain operations
      val = fcn(val)
    self.promise_dict["args"][self.index] = val
    self.promise_dict["ready"] = all(arg is not None for arg in self.promise_dict["args"])
    if self.promise_dict["ready"]:
      self.promise.trigger_completion()
  def exec_bwd(self, val):
    # Brings relevances from the Origin Node to the Arg Node
    # Assumes that bwd_chain has recorded all node bwd closures leading up to an Arg Node
    for fcn in reverse(self.bwd_chain):
      # Iteratively apply the backward chain operations in reverse order from what they were recorded in
      val = fcn(val)
    self.promise_dict["rins"][self.index] = val
class Promise:
  def __init__(self, output_relevance, op, num_branches, parent_promises)
    self.op = op
    self.parents = parent_promises # And the parent connection is from Promise -> Branch
    self.metadata_dict = {
      "rout": output_relevance,
      "ready": False,
      "complete": False,
      "args": [None] * num_branches,
      "rins": [None] * num_branches,
    # metadata_dict is mutable, so all branches can write to it as a shared location,
    #  using their respective index
    self.branches = [ PromiseBranch(self, i) for i in range(num_branches) ]
  def accumulate_rout(self, val):
    # May be overridden based on the Promise operation type, but this is a general default for simplicity
    self.metadata_dict["rout"] += val
  def trigger_completion(self):
    if len(pending_parents := list(filter(lambda parent: not parent.complete, self.parents))):
      # If Promise has any parent Promise Branches that have not been completed yet, try to trigger them by
      #  setting their args with this Promise’s op result
      # This potentially sets off a mutual recursion loop between PromiseBranch.setarg() and
      #  Promise.trigger_completion() calls.
      for parent in pending_parents:
        parent.setarg(self.op(*self.metadata_dict["args"]))
    else:
      # Otherwise, we can proceed with relevance backprop
      # Custom propagation rule for each operation type
      self.relevance = propagate_relevance(type(self.op), output_relevance, self.metadata_dict["args"])
      for i, branch in enumerate(self.branches):
        # Propagates the distributed relevances from Origin Node back to each end-of-Branch Arg Node
        branch.exec_bwd(self.relevance[i])
        for child in branch.children:
          # self.metadata_dict["rins"][i] is updated within branch.exec_bwd()
          child.accumulate_rout(self.metadata_dict["rins"][i])
      self.metadata_dict["complete"] = True
      # Recurse
      for branch in self.branches:
        for child in branch.children:
          child.trigger_completion()
@@ Appendix C Implementation Details
    Backward Pass: Starting from the output, traverse the graph using the Traversal Heuristic (Def 1). A node is traversed when all output relevance is received, or when it is part of a Promise’s resolution path, and marked as propagated when it distributes relevance to its inputs.
@@
Appendix D Promise Examples
D.1 Basic Promise Mechanism
Consider the following subgraph in a computation graph. Suppose nodes A, B, C, D are all nodes which do not store any forward activations, and E is one which does and allows us to retrieve or recompute its forward pass output. Furthermore, suppose that node A is one which requires some forward activation to compute relevance propagation, and nodes B, C, D are all operations whose backward relevance propagations do not require any forward activations.
ABCDEModel forwardDo not store forward activations
Thus, node A is the limiting factor here. Without this forward activation it requires, we cannot continue relevance propagation. But we know there is some node further down the traversal which saves some forward activation(s), and if we traverse to it, we will also know all the operations between it and node A, giving us a way to reconstruct the activation at node A (see Proposition below).
Proposition 1: For any node $v$ in a computation graph, there exists at least one node descendent to $v$ which stores a forward activation, which we call an Arg Node.
Proof.
We prove this using the programming paradigm of PyTorch. We have that every model output is a function of the model inputs and parameters. By the nature of backpropagation, we must compute the gradients of the loss w.r.t. each parameter, but we require the values of the parameters at each layer to compute the gradients w.r.t. the parameters at the preceding layer.
Since every part of the model is a function of the inputs and parameters in some way, their nodes in the computation graph must have out-degree 0. Therefore, as long as the model inputs are set to track gradients, every sink node of the DAG will contain a parameter or input, which are forward activations.
If any node did not eventually lead to such a sink node, that would imply that none of the nodes along this path require gradient tracking. But then this path should not exist in the computation graph at all.
Now, for each node $v$, consider $fwd_{v}$ to be a closure of the node’s forward operation, and $bwd_{v}$ to be a closure of its backward relevance propagation function.
Let $y_{E}$ be the forward output at node E. Then, we can reconstruct the forward input at node A, $x_{A}$, by computing $x_{A}=fwd_{B}(fwd_{C}(fwd_{D}(y_{E})))$.
Let $R_{A}$ be the accumulated relevance at A from its in-neighbours. When we have recomputed $x_{A}$, we can then compute $R_{B\leftarrow A}=bwd_{A}(R_{A},x_{A})$. And similarly, we compute $R_{E\leftarrow D}=bwd_{D}(bwd_{C}(bwd_{B}(R_{B\leftarrow A})))$.
@@ D.1 Basic Promise Mechanism
D.2 Promise Trees and Resolution
@@
And see that the traversal pointer has remained at C, so we simply continue propagating along its path. Also, since $P_{B}$ is Complete, D and E will also be dequeued in coming iterations to continue propagation.
@@
D.3 Solution to Promise Deadlock: Pre-Promises
@@
The parent-child asymmetry creates the desired behavior: $p_{2}$ can reach its Arg Node, forward-chain the activation back through its parent connection to satisfy $p_{1}$’s requirements, but relevance propagation is blocked at $v$ until the traversal heuristic naturally revisits it (when all inputs have landed).

### v3

@@ 6.1 Architecture-Specific Factors on LRP Computational Overhead
Table 2: Architecture-specific values relating to the Promise System’s impact on LRP efficiency. We include a separate row for the finetuned (FT) version of Flan-T5-large, which uses LoRA adapters, changing some of these values drastically. Depth is in terms of either the number of Transformer layers (T) or convolutional layers (C).
@@
6.2 Operation Coverage
By targeting a bounded set of standard tensor operations, our operation-level approach aims for broad coverage. Table A.5 highlights our coverage on a representative subset of architectures, achieving near-perfect node coverage across diverse modalities without model-specific code. We achieve an aggregate 99.99% node coverage (31,464/31,465 nodes) across a broad suite of 15 models spanning vision, language, audio, and state-space architectures. See Appendix G for the full suite of 15 models.
Table 3: DynamicLRP operation coverage for selected architectures.
Model Modality Covered Nodes Unique Ops covered
ViT-b-16 Dosovitskiy et al. (2021) Vision 779/779 16/16
Wav2Vec2-xls-r-300m Babu et al. (2021) Audio 2,021/2,022 18/19
Llama3.2-1B Grattafiori et al. (2024) Language 1,787/1,787 24/24
Mamba-130m Gu and Dao (2024) Language 3,421/3,421 26/26
@@ 7 Discussion
Although current implementation is specific to PyTorch, our principal intellectual contribution is the framework-agnostic Promise-based design. Auto-differentiation engines universally rely on computation graphs to store operation-level data. Since the mathematical requirements for differentiating functions are invariant (the set of “givens” required for a derivative is a property of the function, not the software), the prerequisite data structures are functionally identical across frameworks. Thus, DynamicLRP’s operation-level logic is portable to any system using dynamic computation graphs.
8 conclusion
We have introduced a new algorithm (DynamicLRP) grounded in operations at the primitive tensor level, and a novel mathematical framework (the Promise System) that enables the extension of LRP to arbitrary model architectures. Together, these innovations represent a framework to practically enable the use of LRP for model developers. Looking into existing XAI literature and making use of some theory, we give a sound justification for the choice of this method as a scalable and accurate method for XAI. Empirically, we demonstrate state-of-the-art performance of DynamicLRP both in terms of model attribution faithfulness and in terms of computational overhead. We believe that while further experimental work could be of benefit, the intellectual contributions of this work are substantial and form the groundwork for a promising approach to XAI.
Appendix A LRP Derivation via Deep Taylor Decomposition
LRP redistributes the prediction score (relevance) backwards through a neural network. The challenge of faithful attribution lies in identifying a meaningful distribution rule for the relevance assigned to input $i$ by output neuron $j$, $R_{i\leftarrow j}$. We take advantage of the Deep Taylor Decomposition framework Montavon et al. (2017); Achtibat et al. (2024) to locally linearize and decompose operations into independent contributions, specifically by computing a first-order Taylor expansion of a function $f_{j}(x)$ at a reference point $x_{0}$:
$f_{j}(x)$ $=f_{j}(x_{0})+\sum_{i}\frac{\partial f_{j}}{\partial x_{i}}\Big|_{x_{0}}(x_{i}-x_{0,i})+\mathcal{O}(|x-x_{0}|^{2})$
$=\sum_{i}\frac{\partial f_{j}}{\partial x_{i}}\Big|_{x_{0}}x_{i}+\tilde{b}_{j}$ (5)
where $\tilde{b}_{j}$ collects the constant terms and approximation error. Assuming the relevance $R_{j}^{l}$ is proportional to the function activation $f_{j}(x)$, we can multiply each term in the expansion by a constant ratio $\frac{R_{j}^{l}}{f_{j}(x)}$ to obtain a recursion for relevance:
$R_{j}^{l}=\sum_{i}\underbrace{\frac{\partial f_{j}}{\partial x_{i}}\Big|_{x_{0}}x_{i}\frac{R_{j}^{l}}{f_{j}(x)}}_{R_{i\leftarrow j}}+\underbrace{\tilde{b}_{j}\frac{R_{j}^{l}}{f_{j}(x)}}_{R_{b\leftarrow j}}$
@@ Appendix B Algorithms
Algorithm 3 Computation Graph Pre-Processing
1:Input: Model output hidden_states
2:Initialize $\texttt{in\_adj}\leftarrow\emptyset$, $\texttt{out\_adj}\leftarrow\emptyset$
3:Initialize $\texttt{visited}\leftarrow\emptyset$, $\texttt{topo\_stack}\leftarrow\emptyset$
@@
5:DFS(root, in_adj, out_adj, visited, topo_stack)
6:return in_adj, out_adj, topo_stack $\triangleright$ In-adjacency list, Out-adjacency list, Topologically sorted nodes
7:
8:function DFS(node, in_adj, out_adj, visited, topo_stack)
9:  if node = None or node in visited then $\triangleright$ Skip processed
10:    return
11:  end if
12:  visited.add(node) $\triangleright$ Mark visited
13:  for each child in node.next_functions do
14:    out_adj[node].append(child) $\triangleright$ Record dependencies
15:    in_adj[child].append(node)
16:    DFS(child, in_adj, out_adj, visited, topo_stack) $\triangleright$ Recurse
17:  end for
18:  topo_stack.push(node) $\triangleright$ Add to sort
19:end function
@@
1:Input: Autograd Node node, Propagation input $R_{out}$
2:Output: Propagation output list $R_{in}$
3:if $R_{out}$ is a Promise Branch then $\triangleright$ Record chain if exploring
4:  Define fwd as a closure of node’s forward pass.
5:  Define bwd as a closure of node’s relevance distribution logic.
6:  $R_{out}\texttt{.record(fwd, bwd)}$
7:  return $R_{out}$
8:end if
9:// Propagate $R_{out}$ $\triangleright$ Standard LRP
10:return $R_{in}$
@@
1:Input: Autograd Node node, Propagation input $R_{out}$
2:Output: Propagation output list $R_{in}$
3:if $R_{out}$ is a Promise Branch then $\triangleright$ Retrieve forward activation
4:  Define retrieve_fwd_output as a function that extracts node’s forward pass output given its saved tensors.
5:  activation = retrieve_fwd_output(node)
6:  $R_{out}\texttt{.set\_arg(activation)}$
7:  $R_{out}\texttt{.trigger\_promise\_completion()}$
8:  if $R_{out}\texttt{.promise\_is\_complete}$ then $\triangleright$ Resume propagation
9:    $R_{out}=R_{out}\texttt{.propagated\_relevance}$
10:  else
11:    // Signal for queueing this Node until the Promise is complete
12:    return
13:  end if
14:end if
15:// Propagate $R_{out}$
16:return $R_{in}$
@@
    Backward Pass: Starting from the output, traverse the graph using topological traversal. A node is traversed when all output relevance is received, or when it is part of a Promise’s resolution path, and marked as propagated when it distributes relevance to its inputs.
@@
Appendix C Promise Examples
C.1 Basic Promise Mechanism
In this section, we will provide a complete walk-through for the example provided in Section 4.2. For each node $v$, consider $fwd_{v}$ to be a closure of the node’s forward operation, and $bwd_{v}$ to be a closure of its backward relevance propagation function. Let $y_{E}$ be the forward output at node E. Then, we can reconstruct the forward input at node A, $x_{A}$, by computing $x_{A}=fwd_{B}(fwd_{C}(fwd_{D}(y_{E})))$. Let $R_{A}$ be the accumulated relevance at A from its in-neighbours. When we have recomputed $x_{A}$, we can then compute $R_{B\leftarrow A}=bwd_{A}(R_{A},x_{A})$. And similarly, we compute $R_{E\leftarrow D}=bwd_{D}(bwd_{C}(bwd_{B}(R_{B\leftarrow A})))$.
@@ C.1 Basic Promise Mechanism
C.2 Promise Trees and Resolution
@@
And see that the traversal pointer has remained at C, so we simply continue propagating along its path. Also, since $P_{B}$ is Complete, D and E will also be dequeued in the coming iterations to continue propagation.
@@
C.3 Solution to Promise Deadlock: Pre-Promises
@@
The parent-child asymmetry creates the desired behaviour: $p_{2}$ can reach its Arg Node, forward-chain the activation back through its parent connection to satisfy $p_{1}$’s requirements, but relevance propagation is blocked at $v$ until the traversal heuristic naturally revisits it (when all inputs have landed).
@@
Appendix D Theoretical Analysis
Proposition 1: For any node $v$ in a computation graph, there exists at least one Arg Node which is a descendent of $v$ in the computation graph.
Proof.
Recall from Sec 3.2 that in a computation graph, the sink nodes are the inputs and parameters of the model. To compute the gradients of the loss w.r.t. some parameter at layer $l$, as is the goal of backpropagation, we require the values of the parameters at layer $l+1$ to perform chain rule, assuming the parameters are for linear projection matrices.
Therefore, as long as the model inputs are set to track gradients, every sink node of the DAG will contain a parameter or input, which are forward activations.
If any node did not eventually lead to such a sink node, that would imply that none of the nodes along this path require gradient tracking. But then this path should not exist in the computation graph at all, as it would have no impact on gradient computation.
D.1 DynamicLRP Overhead Complexity
We analyse the computational complexity of our promise-based LRP algorithm in terms of the architectural properties that determine promise resolution requirements.
Definition 3.
$d(v_{p})=\max_{u\in outadj(v_{p})}dist(u,argnode)$
Theorem 1 (Promise-Based LRP Complexity).
Let $\delta=\sum_{v_{p}\in V_{P}}d(v_{p})$
Proof.